OpenStudy (anonymous):

x^2+13x+40 over x-7 divided by x+8 over x^2-49

OpenStudy (anonymous):

\[\Large \frac{ \frac{ x^2+13x+40 }{ x-7 } }{ \frac{ x+8 }{ x^2-49 } }\]

OpenStudy (anonymous):

since you have some thing like that, you would have to know that, \[\Large \frac{ \frac{ a }{ b } }{ \frac{ c }{ d } }~=~\frac{ a }{ b }\times \frac{ d }{ c }\]

OpenStudy (anonymous):

using that, what do you get for the 1st step??

OpenStudy (anonymous):

#13

OpenStudy (anonymous):

Are you there?

OpenStudy (anonymous):

oh....it is the same as what i have written \[\Large \frac{ x^2+13x+40 }{ x-7 }\div \frac{ x+8 }{ x^2-49 }\]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

So whats the 1st step

OpenStudy (anonymous):

okay, so with that said you would have to know that \[\frac{ a }{ b }\div \frac{ c }{ d }=\frac{ a }{ b }\times \frac{ d }{ c }\]

OpenStudy (anonymous):

ok so \[same thing here \div \frac{ x+8 }{ x ^{2}-49 }\]

OpenStudy (anonymous):

oops i meant to flip it lol I'm such a dunce

OpenStudy (anonymous):

you would have to introduce the multiplication sign as you flip it..

OpenStudy (anonymous):

oh yea

OpenStudy (anonymous):

so then what?

OpenStudy (anonymous):

factorise x^2+13x+40

OpenStudy (anonymous):

what do you get??

OpenStudy (anonymous):

http://wolframalpha.com

OpenStudy (anonymous):

what do you get?

OpenStudy (anonymous):

(x+5)(x+8)

OpenStudy (anonymous):

that is right.......are you able to identify x^2-49 as difference of two squares??

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

difference of two squares.. \[a^2-b^2=(a+b)(a-b)\]

OpenStudy (anonymous):

do you know that?

OpenStudy (anonymous):

ok now i do..which #'s are a and b?

OpenStudy (anonymous):

let me use it in an example: \[y^2-4=(y+2)(y-2)\]

OpenStudy (anonymous):

do you get it a bit now??

OpenStudy (anonymous):

oh so what i just did (x+5) and (x+8)

OpenStudy (anonymous):

i do not get what you are saying..

OpenStudy (anonymous):

factorise x^2+13x+40 equals what i just said

OpenStudy (anonymous):

Could you do this part?

OpenStudy (anonymous):

yes.... so if you apply the difference of two square to x^2-49 what do you get?

OpenStudy (anonymous):

do you get it or you are a bit confused?

OpenStudy (anonymous):

(x+5)-(x+8)?

OpenStudy (anonymous):

\[\Large \frac{ (x+5)(x+8) }{ x-7 } \times \frac{ x^2-49 }{ x+8 }\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

you would have to factorise x^2-49 and to do that, difference of two squares is used because both the x^2 and 49 are perfect and squares and are subtracting..

OpenStudy (anonymous):

do you understand it now?

OpenStudy (anonymous):

ok yes...then what would you get?

OpenStudy (anonymous):

i gave an example above try to see whether you can figure it out..

OpenStudy (anonymous):

example: given \[y^2-4\] this could be rewritten as: \[y^2-2^2\] and using the fact i gave for factorising difference of two squares, \[(y+2)(y-2)\]

OpenStudy (anonymous):

ok now i unders

OpenStudy (anonymous):

understand*

OpenStudy (anonymous):

okay.. so what do you get when you do the same for x^2-49??

OpenStudy (anonymous):

\[x ^{2}-7^{2}=(x+7)(x-7)\]

OpenStudy (anonymous):

?hmmm?Was I right?

OpenStudy (anonymous):

yes!....

OpenStudy (anonymous):

so plug in the factorised versions of the x^2+13x+40 and x^2-49 into the question and what do you get?

OpenStudy (anonymous):

ok now can you right out whole problem so i can put it down on paper now that i understand (and maybe a medal too)?

OpenStudy (anonymous):

Oh ok

OpenStudy (anonymous):

i don't get what you are saying...

OpenStudy (anonymous):

x^2+13x+40/x-7 times (x+7)(x-7)/x+8?

OpenStudy (anonymous):

you could just write everything at once \[\Large \frac{ (x+5)(x+8) }{ x-7 }\times \frac{(x+7)(x-7) }{ x+8 }\]

OpenStudy (anonymous):

yay cool

OpenStudy (anonymous):

ok my next question will be the last one like this. :D

OpenStudy (anonymous):

notice some cancel out what do yuo get after cancelling out?

OpenStudy (anonymous):

**you

OpenStudy (anonymous):

Oh...such as?

OpenStudy (anonymous):

(which 1s cancel out there are no common denominators..oh the 7's?

OpenStudy (anonymous):

so x^2 over x+8?

OpenStudy (anonymous):